find the equation of the line that is parallel to the line x=5 and passes throught the point (2,-1)

Answer by Alan3354(69443)

find the equation of the line that is parallel to the line x=5 and passes throught the point (2,-1)
---------------
x=5 means that for any value of, x is 5. That's a line parallel to the y-axis, 5 units to the right of it, a vertical line.
A line parallel to that one, and to the y-axis, thru (2,-1) is x=2.
--------------
Using
y-y1 = m*(x-x1)
y+1 = m(x-2)
But m is undefined, or infinity, so this doesn't help.
You have to figure this one out by logic, or by graphing it.